Find the power output of the engine of a 1200 kg car while the car accelerates from 30 km/h to 100 km/h in 10 s.


1
Expert's answer
2021-09-28T11:12:01-0400

The power is how quickly work can be done. According to kinetic energy theorem, the work is


"W=\Delta E_K=\frac m2(v_2^2-v_1^2)."

Hence, the power is


"P=\frac Wt=\frac m{2t}(v_2^2-v_1^2)=42.13\text{ kW}."